# TIDMAD: Time Series Dataset for Discovering Dark Matter with AI Denoising

### v2

@@ Abstract
Dark matter makes up approximately 85% of total matter in our universe, yet it has never been directly observed in any laboratory on Earth. The origin of dark matter is one of the most important questions in contemporary physics, and a convincing detection of dark matter would be a Nobel-Prize-level breakthrough in fundamental science. The ABRACADABRA experiment was specifically designed to search for dark matter. Although it has not yet made a discovery, ABRACADABRA has produced several dark matter search results widely endorsed by the physics community. The experiment generates ultra-long time-series data at a rate of 10 million samples per second, where the dark matter signal would manifest itself as a sinusoidal oscillation mode within the ultra-long time series. In this paper, we present the TIDMAD — a comprehensive data release from the ABRACADABRA experiment including three key components: an ultra-long time series dataset divided into training, validation, and science subsets; a carefully-designed denoising score for direct model benchmarking; and a complete analysis framework which produces a community-standard dark matter search result suitable for publication as a physics paper. This data release enables core AI algorithms to extract the signal and produce real physics results thereby advancing fundamental science. The data downloading and associated analysis scripts are available at https://github.com/jessicafry/TIDMAD.
@@ 1 Introduction
The quest to uncover the nature of dark matter is one of the biggest challenges in contemporary physics. Several key observations in astrophysics and cosmology have confirmed the existence of dark matter, which constitutes approximately 85% of all mass in the universe [23, 29, 28, 3, 2]. However, dark matter has never been detected by any detector on Earth. Because the composition of dark matter is unknown, theoretical physicists propose various dark matter candidates — hypothetical particles that can be characterized by their physical parameters. Experimental physicists then design experiments to search for these candidates. A convincing detection of any dark matter candidate would be a Nobel-Prize-level breakthrough in fundamental science, but even if nothing is detected, the null results still play a significant role in advancing our understanding of physics by setting limits within the physical parameter space. This means that a particular experiment has eliminated the existence of a dark matter candidates within these limits and does not have sufficient sensitivity to test outside these limits. These limits can be reciprocally used by theoretical physicists to propose better dark matter candidates, thereby improving our understanding of this mysterious constituent of our universe.
@@
Attributable to its extremely rare interactions with normal matter, the signal of dark matter is often submerged in a sea of noise from various sources internal and external to the experimental apparatus. Machine learning (ML) offers a promising means to reduce this noise. Advancements in denoising techniques using ML algorithms have the potential to significantly improve dark matter analyses [24]. These techniques enable the detection of weaker dark matter signals, or in the case of no observation, the setting of stronger limits. In other words, improvements in data denoising directly enhance the scientific reach of dark matter experiments. In this paper, we present an ultra-long time series dataset produced by a real dark matter detector: ABRACADABRA (A Broadband/Resonant Approach to Cosmic Axion Detection with an Amplifying B-field Ring Apparatus, abbrev. ABRA-10cm). ABRA-10cm is the world leading sub-$\mu eV$ dark matter experiment that pioneered the quantum enabled lumped element dark matter detection technique [19, 20, 25]. We operated the ABRA-10cm detector in February 2024 to obtain a special time series dataset for these studies: TIDMAD (TIme series dataset for discovering Dark Matter with Ai Denoising). It can be partitioned into three parts: (1) training data, (2) validation data, and (3) science data.
@@
The training data include time series data where a dark matter-like signal is injected by hardware. If dark matter enters ABRA, it will manifest itself as a sinusoidal oscillation mode within the time series; therefore, the injected signals are also sinusoidal oscillations at specific frequencies. Both the detected (noisy) time series and the injected (ground truth, clean) time series are provided with one-to-one temporal correspondence. This allows the training of machine learning algorithms to denoise the detector data and recover the injected signal. The validation data is used to produce Benchmark 1: Denoising Score, see Section 4.1. Algorithms that effectively dampen the detector noise while amplifying the injected signal will achieve a better denoising score. The science data is collected without the injected signal with an extended duration to produce Benchmark 2: Dark Matter Limits, as discussed in the previous paragraph. The limit generation procedure is detailed in Section 4.2. The scientific data are titled to reflect their use in producing real, community-standard physics results that are suitable for presentation in scientific journals. Several traditional and deep learning denoising algorithms are presented in Section 3 and Appendix C, where the resulting denoised data is benchmarked against the raw, un-denoised detector data.
@@ 1.1 Axion dark matter and ABRACADABRA
In recent years, the axion has emerged as one of the leading dark matter candidates as a result of its theoretical elegance. Axions interact with normal matter via electromagnetism, which can be characterized by a physics parameter $g_{a\gamma\gamma}$. Arising from its small mass $m_{a}<1eV$ ($10^{-6}$ times smaller than electron), axions acts as a classical field oscillating at a frequency $f_{a}=m_{a}/2\pi$. Astrophysical measurements determine that the Earth exists in a bath of dark matter with a known local density of $\rho_{DM}$ [11].
Figure 1: Left to right: ABRA-10cm dilution fridge with outer vacuum cans on; Coldest stage of ABRA-10cm fridge above shielded 1T superconducting torroidal magnet; Interior of ABRA-10cm magnet including pickup and calibration loop wires existing the center of the magnet; Effective circuit diagram for both dark matter and injected signals.
The latest advancements in quantum detector technology have facilitated new avenues to search for the axion. ABRA-10cm is one of the novel detectors designed to search for axions leveraging these advancements in quantum technologies [20]. ABRA-10cm capitalizes on the fact that we are immersed in a bath of a feebly electromagnetically-interacting, oscillating dark matter field to detect this elusive particle. Specifically, in the presences of a static magnetic field $\textbf{B}_{0}$, the axion, henceforth referred to as dark matter, induces an oscillating magnetic field $\textbf{B}_{a}$. Thus, to detect dark matter ABRA-10cm provides a strong magnetic field $B_{0}=1T$ and uses a superconducting pickup loop to read out the oscillating dark matter signal. Read out by a superconducting quantum interference device (SQUID), the pickup loop detects the dark matter signal as a time-oscillating current given by
@@ 2 Dataset description
The data presented in this paper were acquired using the ABRACADABRA detector. The overall schematics of the data is shown in Figure LABEL:fig:dataflow. All data are saved as a series at $10$ MS/s (Megasample per second), where each sample is a 8-bit integer ranging from -128 to 127. These integers can be converted into a physics units of mV (millivolts) with a scaling factor of $40/128$. The procured datasets are stored at Open Science Data Federation (OSDF) [34] in .hdf5 format and can be accessed via the download_data.py script in the github repository provided.
Figure 3: TIDMAD data flow explaining how data splits (green) result in benchmarks (blue). Rectangles correspond to provided scripts. Left to right, top to bottom the scripts are , , , , fig:dataflow
@@ 3 Experiments
We experimented five different denoising algorithms including two traditional algorithms and three deep learning models. The traditional algorithm can be directly applied to the validation dataset using inference.py, while deep learning models need to be trained first on the training dataset with train.py. The five algorithms are listed below:
@@
  • Savitzky-Golay filter: with a window size of 19 and a polynomial order of 11
  • Fourier Averaging (baseline): This operation was originally adopted by the ABRACADABRA experiment to compress its data [19, 25]. For every 10-second time series segment, a Fourier transform is applied to each one-second interval, and the resulting 10 Fourier spectra are averaged to produce a single spectrum representing the 10-second segment. This method is considered the baseline approach, representing the current status quo of ABRACADABRA.
  • FC net: an autoencoder architecture designed for transforming input data. This model consists of an encoder and a decoder. The encoder encodes the input data into a low-dimensional representation, while the decoder reconstructs the original data from this encoded representation. Both the encoder and decoder are composed of multiple fully-connected layers and activation layers. FC-Net outputs a single float point number at each time step, and the training is conducted by minimizing the mean square error between this float point number and corresponding ground truth time series at every sample.
  • WaveNet [30]: a deep neural network that was originally developed to generates high-quality raw audio waveforms by autoregressively predicting each sample using dilated causal convolutions. The model was adapted with residual blocks containing exponentially increasing dilation rates, gated activations, and skip connections to maintain the same input-output shape for TIDMAD denoising task.
@@
  • RNN Sequence to Sequence Model [30]: a neural architecture with separate encoder and decoder RNNs where the encoder processes the entire input sequence to produce a context representation, and the decoder autoregressively generates the output sequence. Both encoder and decoder adopts a LSTM architecture, and the decoder output is the same 256-class classification decision as PU-Net. PU-Net, Transformer, WaveNet, and RNN Seq2Seq are all trained using Focal Loss to handle class-imbalanced segmentation labels  [16].
@@ 4.1 Benchmark 1: denoising score
The denoising score is a modified signal-to-noise ratio (SNR) of the denoised CH1 SQUID time series. It is calculated over the validation dataset by first segmenting both the injected and SQUID time series into one-second segments. Each second of the time series is transformed into a power spectral density (PSD) using a squared fast Fourier transform. This frequency domain data records signal power at each frequency -- $PSD(\nu)$. Since the injected dark matter signal is a clean sinusodial oscillation, it should appear as a single-bin peak ($\nu_{0}$) in the PSD, while noise in SQUID time series is distributed across all frequency bins. The location of $\nu_{0}$ is identified by the PSD of the injected time series (ground truth) as the largest single bin peak relative to its nearest neighbors.
@@
where $df$ is the sampling frequency of $10^{-7}$ Hz. Once $\nu_{0}$ is identified in the injected time series, the the signal region is defined by selecting $n_{sig}=\pm 1$ bins around the signal frequency to account for spectral leakage. Similarly, the noise region is defined by selecting $n_{bkgd}=\pm 50$ bins outside of the signal region. By taking the ratio of the PSD in the signal region to that in the noise region, we acquire the SNR for each one-second segment PSD.
@@
The SQUID SNR then gets multiplied to the corresponding, normalized SQUID SNR in the the same one-second segments, and then summed over all one-second segments to produce $\Lambda$ defined below:
@@
We examine the validity of $\Lambda$ to represent denoising efficiency through a study involving added Gaussian noise. In the ABRA-10cm detector, there are a multitude of independent, random noise sources internal and external to the detector thus the noise in our data follows a Gaussin distribution. In this study, Gaussian noise is imposed on the injected time series, and $\Lambda$ are calculated at different levels of Gaussian noise amplitudes and standard deviations. As shown in Figure 4 (right), $\Lambda$ exhibits an exponential decay trend with increased Gaussian noise amplitude. To establish a linear correlation between the denoising score and noise, we apply a logarithmic transformation to $\Lambda$ to calculate the final denoising score:
@@
Fourier Averaging $1\times 10^{8}$ N/A 10-fold Average 0.24 0.26
Moving Average $1\times 10^{6}$ N/A window = 100 0.86 0.95
SG Filter $1\times 10^{6}$ N/A window = 19, order = 11 0.95 1.04
@@
WaveNet $4\times 10^{4}$ N/A See Appendix  B 4.99 5.16
@@
Table 1 shows the denoising scores for all algorithms discussed in Section 3. The case with no denoising is shown in the first row, with its fine denoising score calibrated to 1. We observed that in all cases, the coarse denoising score is slightly higher than the fine denoising score. Based on the results, all traditional algorithms decrease the denoising score because time domain averaging erases high-frequency signals in the region of interest. Meanwhile, all deep learning algorithms efficiently boost the denoising score. Surprisingly, we observed that the FC Net model achieved the best performance with a denoising score of 6.43.
@@ 4.2 Benchmark 2: dark matter limit
The second benchmark empowers algorithm creators with the capability to conduct a community-standard dark matter search using the science dataset. Following the discussion in the Introduction, the two physics parameters for the dark matter candidate in this paper are the dark matter mass ($m_{a}$) and the dark matter to electromagnetic coupling ($g_{a\gamma\gamma}$). Null results from different dark matter experiments place limits within this parameter space expressed by the shaded regions in Figure 4.2. In the physics community, a better dark matter limit is represented by pushing towards lower values of $g_{a\gamma\gamma}$ at different $m_{a}$. We provide a comprehensive tool necessary for performing the statistical analysis to produce dark matter limits in Figure 4.2. The dark matter mass $m_{a}$ is directly proportional to frequency, therefore the limit-setting procedure is repeated for 11.1 million independent $m_{a}$ from 0.4 neV (100 kHz) to 8 neV (2 MHz). The dark matter limit at each $m_{a}$ is obtained using a frequentist log-likelihood ratio test statistic (TS), with the results depicted as regions in Figure 4.2. The detail of this analysis can be found in Appendix  C.
@@
Figure 5: Plotted alongside the present, state-of-the-art axion dark matter limits (grey) are the $2\sigma$ exclusion limits for ABRA-10cm Run 3 (pink) [25], ABRA-TIDMAD Raw limit (red), and ABRA-TIDMAD Denoised limit from the trained FC net (maroon). Shaded regions correspond to pairing of dark matter model parameters ($m_{a}$, $g_{a\gamma\gamma}$) that are ruled out by the specified experiment or observation, and the bright yellow region indicates theoretical predictions (See Appendix D for details). The plotting script is modified from the publicly available repository [17].

### v3

@@ Abstract
Dark matter makes up approximately 85% of total matter in our universe, yet it has never been directly observed in any laboratory on Earth. The origin of dark matter is one of the most important questions in contemporary physics, and a convincing detection of dark matter would be a Nobel-Prize-level breakthrough in fundamental science. The ABRACADABRA experiment was specifically designed to search for dark matter. Although it has not yet made a discovery, ABRACADABRA has produced several dark matter search results widely endorsed by the physics community. The experiment generates ultra-long time-series data at a rate of 10 million samples per second, where the dark matter signal would manifest itself as a sinusoidal oscillation mode within the ultra-long time series. In this paper, we present the TIDMAD — a comprehensive data release from the ABRACADABRA experiment including three key components: an ultra-long time series dataset divided into training, validation, and science subsets; a carefully-designed denoising score for direct model benchmarking; and a complete analysis framework which produces a physics community-standard dark matter search result suitable for publication as a physics paper. This data release enables core AI algorithms to extract the dark matter signal and produce real physics results thereby advancing fundamental science. The data downloading and associated analysis scripts are available at https://github.com/jessicafry/TIDMAD.
@@ 1 Introduction
The quest to uncover the nature of dark matter is one of the biggest challenges in contemporary physics. Several key observations in astrophysics and cosmology have confirmed the existence of dark matter, which constitutes approximately 85% of all mass in the universe [26, 32, 31, 3, 2]. However, dark matter has never been detected by any detector on Earth. Because the composition of dark matter is unknown, theoretical physicists propose various dark matter candidates — hypothetical particles that can be characterized by their physical parameters. Experimental physicists then design experiments to search for these candidates. A convincing detection of any dark matter candidate would be a Nobel-Prize-level breakthrough in fundamental science, but even if nothing is detected, the null results still play a significant role in advancing our understanding of physics by setting limits within the physical parameter space. This means that a particular experiment has eliminated the existence of a dark matter candidate within these limits and does not have sufficient sensitivity to test outside these limits. Reciprocally, these limits are used by theorists to propose better dark matter candidates, thereby improving our understanding of this mysterious constituent of our universe.
@@
Attributable to its extremely rare interactions with normal matter, the signal of dark matter is often submerged in a sea of noise from various sources internal and external to the experimental apparatus. Machine learning (ML) offers a promising means to reduce this noise. Advancements in denoising techniques using ML algorithms have the potential to significantly improve dark matter analyses [27]. These techniques enable the detection of weaker dark matter signals, or in the case of no observation, the setting of stronger limits. In other words, improvements in data denoising directly enhance the scientific reach of dark matter experiments. In this paper, we present an ultra-long time series dataset produced by a real dark matter detector: ABRACADABRA (A Broadband/Resonant Approach to Cosmic Axion Detection with an Amplifying B-field Ring Apparatus, abbrev. ABRA-10cm). ABRA-10cm is the world leading sub-$\mu eV$ dark matter experiment that pioneered the quantum enabled lumped element dark matter detection technique [21, 22, 28]. We operated the ABRA-10cm detector in February 2024 to obtain a special time series dataset for these studies: TIDMAD (TIme series dataset for discovering Dark Matter with Ai Denoising). These data are partitioned into three parts: (1) training data, (2) validation data, and (3) science data.
@@
The training data include time series data where a dark matter-like signal is injected by hardware. If dark matter enters ABRA-10cm, it will manifest itself as a sinusoidal oscillation mode within the time series; therefore, the injected signals are also sinusoidal oscillations within a range of specific frequencies. While the signal shape is known, the dark matter signal amplitude and frequency are unknown parameters in dark matter searches. Both the detected (noisy) time series and the injected (ground truth, clean) time series are provided with one-to-one temporal correspondence. This allows the training of machine learning algorithms to denoise the detector data and recover the injected signal. The validation data is used to produce Benchmark 1: Denoising Score, see Section 4.1. Algorithms that effectively dampen the detector noise while amplifying the injected signal will achieve a better denoising score. The science data is collected without the injected signal with an extended duration to produce Benchmark 2: Dark Matter Limits. The limit generation procedure is detailed in Section 4.2. The scientific data are titled to reflect their use in producing real, community-standard physics results that are suitable for presentation in scientific journals. Several traditional and deep learning denoising algorithms are presented in Section 3 and Appendix C, where the resulting denoised data is benchmarked against the raw, un-denoised detector data.
@@ 1.1 Axion dark matter and ABRACADABRA
In recent years, the axion has emerged as one of the leading dark matter candidates as a result of its theoretical elegance. Axions interact with normal matter via electromagnetism, which can be characterized by a physics parameter $g_{a\gamma\gamma}$. Arising from its small mass $m_{a}<1eV$ ($10^{-6}$ times smaller than electron), axions act as a classical field oscillating at a frequency $f_{a}=m_{a}/2\pi$. Astrophysical measurements determine that the Earth exists in a bath of dark matter with a known local density of $\rho_{DM}$ [12].
Figure 1: Left to right: ABRA-10cm dilution fridge with outer vacuum cans on; Coldest stage of ABRA-10cm fridge above shielded 1T superconducting torroidal magnet; Interior of ABRA-10cm magnet including pickup and calibration loop wires in the center of the magnet; Effective circuit diagram for both dark matter and injected signals.
The latest advancements in quantum detector technology have facilitated new avenues to search for the axion. ABRA-10cm is one of the novel detectors designed to search for axions leveraging these advancements in quantum technologies [22]. ABRA-10cm capitalizes on the fact that we are immersed in a bath of a feebly electromagnetically-interacting, oscillating dark matter field to detect this elusive particle. Specifically, in the presence of a static magnetic field $\textbf{B}_{0}$, the axion, henceforth referred to as dark matter, induces an oscillating magnetic field $\textbf{B}_{a}$. Thus, to detect dark matter ABRA-10cm provides a strong magnetic field $B_{0}=1T$ and uses a superconducting pickup loop to observe the oscillating dark matter signal. Read out by a superconducting quantum interference device (SQUID), the pickup loop detects the dark matter signal as a time-oscillating current given by
@@
The frequency of this oscillating signal is a model parameter meaning if we knew the dark matter mass, this frequency would be set. However, theoretical models point to a range of possible dark matter masses, not a singular value; this signal frequency range covers more than seven orders of magnitude making the data denoising task extremely challenging.
@@ 2 Dataset description
The data presented in this paper were acquired using the ABRACADABRA detector. The overall schematics of the data is shown in Figure 3. All data are saved as a series at $10$ MS/s (Megasample per second), where each sample is a 8-bit integer ranging from -128 to 127. These integers can be converted into a physics units of mV (millivolts) with a scaling factor of $40/128$. The procured datasets are stored at Open Science Data Federation (OSDF) [37, 6, 18] in .hdf5 format and can be accessed via the download_data.py script in the github repository provided.
Figure 3: TIDMAD data flow explaining how data splits (green) result in benchmarks (blue). Rectangles correspond to provided scripts. Left to right, top to bottom the scripts are train.py, inference.py, benchmark.py, process_science_data.py, brazilband.py
@@ 3 Experiments
We benchmarked eight different denoising algorithms including three traditional algorithms and five deep learning models. The traditional algorithm can be directly applied to the validation dataset using inference.py, while deep learning models need to be trained first on the training dataset with train.py. The eight algorithms are listed below:
@@
  • Savitzky-Golay filter: with a window size of 19 and a polynomial order of 11.
  • Fourier Averaging (baseline): this operation was originally adopted on ABRA-10cm to compress its data [21, 28]. For every 10-second time series segment, a Fourier transform is applied to each one-second interval, and the resulting 10 Fourier spectra are averaged to produce a single spectrum representing the 10-second segment. This method is considered the baseline approach, representing the current status quo of ABRA-10cm.
  • FC net: an autoencoder architecture designed for transforming input data. This model consists of an encoder and a decoder. The encoder encodes the input data into a low-dimensional representation, while the decoder reconstructs the original data from this encoded representation. Both the encoder and decoder are composed of multiple fully-connected layers and activation layers. FC-Net outputs a single floating point number at each time step, and the training is conducted by minimizing the mean square error between this floating point number and corresponding ground truth time series at every sample.
  • WaveNet: a deep neural network originally developed to generates high-quality raw audio waveforms by autoregressively predicting each sample using dilated causal convolutions [33]. The model was adapted with residual blocks containing exponentially increasing dilation rates, gated activations, and skip connections to maintain the same input-output shape for TIDMAD denoising task.
@@
  • RNN Sequence to Sequence Model: a neural architecture with separate encoder and decoder RNNs where the encoder processes the entire input sequence to produce a context representation, and the decoder autoregressively generates the output sequence  [33]. Both encoder and decoder adopts a LSTM architecture, and the decoder output is the same 256-class classification decision as PU-Net. PU-Net, Transformer, WaveNet, and RNN Seq2Seq are all trained using Focal Loss to handle class-imbalanced segmentation labels  [17].
@@ 4.1 Benchmark 1: denoising score
The denoising score is a modified signal-to-noise ratio (SNR) of the denoised CH1 SQUID time series. It is calculated over the validation dataset by first segmenting both the injected and SQUID time series into one-second segments. Each second of the time series is transformed into a power spectral density (PSD) using a squared fast Fourier transform. This frequency domain data records signal power at each frequency – $PSD(\nu)$. Since the injected dark matter signal is a clean sinusodial oscillation, it should appear as a single-bin peak ($\nu_{0}$) in the PSD, while noise in SQUID time series is distributed across all frequency bins. The location of $\nu_{0}$ is identified by the PSD of the injected time series (ground truth) as the largest single bin peak relative to its nearest neighbors.
@@
where $df$ is the sampling frequency of $10^{-7}$ Hz. Once $\nu_{0}$ is identified in the injected time series, the signal region is defined by selecting $n_{sig}=\pm 1$ bins around the signal frequency to account for spectral leakage. Similarly, the noise region is defined by selecting $n_{bkgd}=\pm 50$ bins outside of the signal region. By taking the ratio of the PSD in the signal region to that in the noise region, we acquire the SNR for each one-second segment PSD.
@@
The SQUID SNR then gets multiplied to the corresponding, normalized SQUID SNR in the same one-second segments, and then summed over all one-second segments to produce $\Lambda$ defined below:
@@
We examine the validity of $\Lambda$ as a measure of denoising efficiency through a controlled study involving added Gaussian noise. The ABRA-10cm detector is subject to a variety of independent noise sources both internal and external, and both stable and stochastic in time. Together, these real noise sources produce a distinctly non-Gaussian noise distribution, as demonstrated in Figure 2. Although the detector noise itself is non-Gaussian, we introduce synthetic Gaussian noise to the real time series data to probe how $\Lambda$ behaves under systematically varied and well-characterized noise conditions. This approach allows the denoising score to be linearized and validated in a controlled setting, even if not precisely replicating the complex, full statistical structure of the detector noise. In this study, Gaussian noise is added to the SQUID time series and $\Lambda$ values are computed for a range of added noise amplitudes and standard deviations. As shown in Figure 4 (right), $\Lambda$ exhibits an exponential decay trend with increased noise amplitude. To establish a linear correlation between the denoising score and noise, we apply a logarithmic transformation to $\Lambda$ to obtain the final denoising score:
@@
Fourier Averaging $1\times 10^{8}$ 10-fold Average 0.24 0.26
Moving Average $1\times 10^{6}$ window = 100 0.86 0.95
SG Filter $1\times 10^{6}$ window = 19, order = 11 0.95 1.04
@@
WaveNet $4\times 10^{4}$ No See Appendix  B 4.99 5.16
@@
Table 1 shows the denoising scores for all algorithms discussed in Section 3. The case with no denoising is shown in the first row, with its fine denoising score calibrated to 1. The deep learning algorithms repeatedly outperform all of the traditional denoising methods. Based on the results, all traditional algorithms decrease the denoising score because time domain averaging erases high-frequency signals in the region of interest. Meanwhile, all deep learning algorithms efficiently boost the denoising score. We observed that in all cases, the coarse denoising score is slightly higher than the fine denoising score. Surprisingly, we observed that the FC Net model achieved the best performance with a denoising score of 6.43.
@@ 4.2 Benchmark 2: dark matter limit
The second benchmark empowers algorithm creators with the capability to conduct a community-standard dark matter search using the science dataset. This benchmark bridges the gap between ML and particle physics; ML developers can translate algorithmic performance directly to improved particle physics experimental reach. By abstracting the full physics analysis chain, the ML community can translate their denoising results to physics results without understanding the underlying physics, all while maintaining full physics rigor.
The two physics parameters for the dark matter candidate in this paper are the dark matter mass ($m_{a}$) and the dark matter to electromagnetic coupling ($g_{a\gamma\gamma}$). Null results from different dark matter experiments place limits within this parameter space expressed by the shaded regions in Figure 5. In the physics community, a better dark matter limit is represented by pushing towards lower values of $g_{a\gamma\gamma}$ at different $m_{a}$. For dark matter physicists, this dark matter exclusion limit is the community standard for benchmarking different detector performances against each other on the metric of experimental sensitivity to theoretical dark matter candidates. We provide a comprehensive tool necessary for performing the statistical analysis to produce dark matter limits in Figure 5. The limit-setting procedure is repeated for 11.1 million independent $m_{a}$ ($f_{a}$) from 0.4 neV (100 kHz) to 8 neV (2 MHz). The dark matter limit at each $m_{a}$ is obtained using a frequentist log-likelihood ratio test statistic (TS). The details of this analysis can be found in Appendix  C. However, by simply inputting the provided science data, denoised with an algorithm of choice, into the brazilband.py script, ML developers can automatically produce Figure 5 to measure the improvements to physics experimental reach enabled by their denoising algorithm.
@@
Figure 5: Plotted alongside the present, state-of-the-art axion dark matter limits (grey) are the $2\sigma$ exclusion limits for ABRA-10cm Run 3 (pink) [28], ABRA-TIDMAD Raw limit (red), and ABRA-TIDMAD Denoised limit from the trained FC net (maroon). The ABRA-TIDMAD result does not surpass ABRA-10cm Run 3 due to reduced geometric coupling and shorter data taking duration in the TIDMAD configuration. Shaded regions correspond to pairings of dark matter model parameters ($m_{a}$, $g_{a\gamma\gamma}$) that are ruled out by the specified experiment, and the bright yellow region indicates theoretical predictions (See Appendix D for details). The plotting script is modified from the publicly available AxionLimits repository [19].